# REDUCIO! Generating 1024$\times$1024 Video within 16 Seconds using Extremely Compressed Motion Latents

### v1

@@ 4.1 Training and Evaluation Details
Our Reducio-VAE is trained on a 400K video dataset collected from Pexels***https://www.pexels.com/. Pexels contains a large number of free and high-quality stock videos, where each video is accompanied with a short text description. We train Reducio-VAE from scratch on $256\times 256$ videos with 16 FPS. Note that we split videos with higher resolutions into tiles in the spatial dimension. To perform the VAE comparison, PSNR, SSIM [50], LPIPS [63] and reconstruction FVD (rFVD) [45] are employed as the evaluation metrics.
@@ 4.2 Main Results
As shown in the table, Reducio-VAE strikes obvious advantages on overall reconstruction performance. Specifically, our model significantly outperforms state-of-the-art 2D VAE, e.g., SD2.1-VAE and SDXL-VAE, by more than 5 db in PSNR. Moreover, Reducio-VAE also demonstrates better performance compared with VAEs designed for video in recent literature, e.g., OmniTokenizer and OpenSora-1.2. Our VAE also outperforms the concurrent work, Cosmos-VAE, which also adopts a further compressed video latent space, by 0.2 in SSIM and 5 db in PSNR with an 8×8$\times$8 × higher down-sampling factor.
We observe Reducio-VAE has worse rFVD than others on UCF-101. By careful inspection, we find the selected prior content image is blurring in many cases due to the low visual quality of UCF-101 and hence result in a lower recontsruction quality. In contrast, Reducio-VAE places first on rFVD on pexels, which consists videos with much improved viusal quality and much higher resolution.
@@
Moreover, Reducio-DiT significantly boost the throughput of video LDMs. The compressed latent space facilitates Reducio-DiT to output $256^{2}$ videos of 16 frames in almost 1 second, which realize a $\mathbf{5.3\times}$ speed up compared with Lavie. Moreover, 16-frame $1024^{2}$ video generation costs Reducio-DiT only for 15.4 seconds on average, leading to a striking speedup of $\mathbf{16.6\times}$ over Lavie. We argue the core philosophy of Reducio lies in the extremely compressed latent space, which acts as the key role in accelerating both training and inference speed of video LDMs.
@@ 4.3.1 Design of Reducio-VAE
Fusion with cross-attention achieves the most competitive performance in Tab. 5. As for add-based fusion, we feed the content frame features into a convolution and replicate the output on temporal dimension. Then we add video features with the content feature, together with a temporal embedding. Concerning linear-based fusion, we view the content condition as a single-frame video feature and concatenate it with the video feature by the temporal dimension. The concatenated features are then refined by a 3D convolution layer. As observed in Tab. 5, the attention-based Reducio-VAE outperforms other baselines by 1.71 in PSNR and 0.01 in SSIM, while demanding higher computational costs. By default, we display results using the attention-based fusion. Note that all Reducio-VAE displayed in Tab. 5 share the same 3D encoder and hence an identical latent space, where users are free to choose any of the decoders with different resource constraints.

### v2

@@ 4.1 Training and Evaluation Details
Our Reducio-VAE is trained on a 400K video dataset collected from Pexels***https://www.pexels.com/. Pexels contains a large number of free and high-quality stock videos, where each video is accompanied by a short text description. We train Reducio-VAE from scratch on $256\times 256$ videos with 16 FPS. Note that we split videos with higher resolutions into tiles in the spatial dimension. To perform the VAE comparison, PSNR, SSIM [54], LPIPS [67] and reconstruction FVD (rFVD) [49] are employed as the evaluation metrics.
@@ 4.2 Main Results
As shown in the table, Reducio-VAE strikes obvious advantages on overall reconstruction performance. Specifically, our model significantly outperforms state-of-the-art 2D VAE, e.g., SD2.1-VAE and SDXL-VAE, by more than 5 db in PSNR. Moreover, Reducio-VAE also performs better than VAEs designed for video in recent literature, e.g., OmniTokenizer and OpenSora-1.2. Our VAE also outperforms the concurrent work, Cosmos-VAE, which also adopts a further compressed video latent space, by 0.2 in SSIM and 5 db in PSNR with an 8×8$\times$8 × higher down-sampling factor.
We observe Reducio-VAE has worse rFVD than others on UCF-101. By careful inspection, we find the selected prior content image is blurring in many cases due to the low visual quality of UCF-101 and hence result in a lower reconstruction quality. In contrast, Reducio-VAE places first on rFVD on pexels, which consists of videos with much improved visual quality and much higher resolution.
@@
Moreover, Reducio-DiT significantly boosts the throughput of video LDMs. The compressed latent space facilitates Reducio-DiT to output $256^{2}$ videos of 16 frames in almost 1 second, which realize a $\mathbf{5.3\times}$ speed up compared with Lavie. Moreover, 16-frame $1024^{2}$ video generation costs Reducio-DiT only for 15.4 seconds on average, leading to a striking speedup of $\mathbf{16.6\times}$ over Lavie. We argue the core philosophy of Reducio lies in the extremely compressed latent space, which acts as the key role in accelerating both the training and inference speed of video LDMs.
@@ 4.3.1 Design of Reducio-VAE
Fusion with cross-attention achieves the most competitive performance in Tab. 5. As for add-based fusion, we feed the content frame features into a convolution and replicate the output on the temporal dimension. Then we add video features with the content feature, together with a temporal embedding. Concerning linear-based fusion, we view the content condition as a single-frame video feature and concatenate it with the video feature by the temporal dimension. A 3D convolution layer then refines the concatenated features. As observed in Tab. 5, the attention-based Reducio-VAE outperforms other baselines by 1.71 in PSNR and 0.01 in SSIM, while demanding higher computational costs. By default, we display results using the attention-based fusion. Note that all Reducio-VAE displayed in Tab. 5 share the same 3D encoder and hence an identical latent space, where users are free to choose any of the decoders with different resource constraints.
@@ 5 Conclusion
Reducio is purely a research project. Currently, we have no plans to incorporate Reducio into a product or expand access to the public. We will also put Microsoft AI principles into practice when further developing the models. In our research paper, we account for the ethical concerns associated with text-image-to-video research. To mitigate issues associated with training data, we have implemented a rigorous filtering process to purge our training data of inappropriate content, such as explicit imagery and offensive language, to minimize the likelihood of generating inappropriate content.
Appendix A Implementation details
Reducio-VAE. We demonstrate the detailed architecture of Reducio in Fig. 8. Generally, we follow the training strategies of SD-VAE [40]. We employ a tailored version of PatchGAN [18] based on 3D convolutions and optimize the model with $L_{1}$ loss, KL loss, perceptual loss [68], and GAN loss. While initializing the 2D encoder and the 3D VAE with SD-VAE pre-trained weights accelerates convergence, we find that freezing the 2D encoder leads to worse performance than training the full parameters. We follow LaMD [16] to feed the motion latent into a normalization layer to obtain the output of the VAE encoder. Therefore, in the stage of diffusion training, we use a scale factor of 1.0, which is multiplied by the input latent as inputs into DiT.
Figure 8: The detailed architecture of Reducio-VAE with $f_{t}=4,f_{s}=32$.
Table 9: Hyperparameters for Reducio-VAE
Reducio-VAE
$z$-shapeon $256^{2}$ videos $4\times 8\times 8\times 16$ $8\times 8\times 8\times 16$
Channels (3d) 128
Channels (2d) 128
Ch Multiplier (3d) 1,2,2,4,4,4
Ch Multiplier (2d) 1,2,2,4,4
Depth 2
Batch size 32 24
Learning rate 4e-5
Iterations 1,000,000
Table 10: Comparison with more SOTA models on Vbench.
Model Quality Semantic Total
Score Score Score
Show-1 [66] 80.42 72.98 78.93
Lavie [53] 78.78 70.31 77.08
VideoCrafter [5] 81.59 72.22 79.72
OpenSora v1.2 [30] 81.35 73.39 79.76
Lavie-2 [53] 83.24 75.76 81.75
Pyramid Flow [19] 84.74 69.62 81.72
VideoCrafter-2 [12] 83.27 76.73 81.97
Reducio-DiT 82.24 78.00 81.39
During inference, We split videos with a resolution over $256\times 256$ into overlapping spatial tiles, we fuse the encoded latent as well as the video output, in a similar manner with Movie Gen [35]. Note that since Reducio-VAE employs a spatial down-sampling factor of 32, Reducio-VAE can only process video inputs whose width and height are divisible by 32. Otherwise, videos should be padded to meet this requirement before being fed into the VAE.
Reducio-DiT. We elaborate on the details of content image conditions in Fig. 9. During inference, we use classifier-free guidance [13] for better generation quality and set the default scale to 2.5. During training, we randomly drop image conditions at a probability of 0.1, as well as drop 10% text conditions.
Figure 9: The overview of the efficient content condition solution for Reducio-DiT on high-resolution videos.
Table 11: Detailed quantitative comparison with SOTA for text-to-video generation models on VBench.
Model subject background temporal motion dynamic aesthetic imaging object multiple human color spatial scene appearance temporal overall
consistency consistency flickering smoothness degree quality quality class objects action relationship style style consistency
Lavie [53] 91.41 97.47 98.30 96.38 49.72 54.94 61.90 91.82 33.32 96.80 86.39 34.09 52.69 23.56 25.93 26.41
Show-1 [66] 95.53 98.02 99.12 98.24 44.44 57.35 58.66 93.07 45.47 95.60 86.35 53.50 47.03 23.06 25.28 27.46
VideoCrafter [5] 95.10 98.04 98.93 95.67 55.00 62.67 65.46 78.18 45.66 91.60 93.32 58.86 43.75 24.41 25.54 26.76
OpenSora v1.2 [30] 96.75 97.61 99.53 98.50 42.39 56.85 63.34 82.22 51.83 91.20 90.08 68.56 42.44 23.95 24.54 26.85
Lavie-2 [53] 97.90 98.45 98.76 98.42 31.11 67.62 70.39 97.52 64.88 96.40 91.65 38.68 49.59 25.09 25.24 27.39
Pyramid Flow [19] 96.95 98.06 99.49 99.12 64.63 63.26 65.01 86.67 50.71 85.60 82.87 59.53 43.20 20.91 23.09 26.23
VideoCrafter-2 [12] 97.17 98.54 98.46 97.75 42.50 65.89 70.45 93.39 49.83 95.00 94.41 64.88 51.82 24.32 25.17 27.57
Reducio-DiT 98.05 99.13 98.45 98.77 27.78 64.02 67.67 91.49 69.91 92.60 89.06 52.85 54.90 25.16 26.40 28.87
Appendix B Additional results
We display the detailed performance comparison on VBench [17] in Tab. 11 and Tab. 10. As the results show, Reducio-DiT achieves a promising semantic score of 78.00, beating a range of state-of-the-art LDMs.